An electrical motor spins at a constant 2695.0 rpm. If the rotor radius is 7.165 cm, what is the linear acceleration of the edge of the rotor?
Answer #1

The question can be solved as follows:

Angular-speed, w = 2695 rpm = 2695*2pi/60 = 282.22 rad/sec,

Radius, r = 7.165 cm = 0.07165 meters,

So, linear-velocity, v = rw = 0.07165*282.22 = 20.22 m/sec,

Hence,

Acceleration, a = v^2/r = (20.22)^2/0.07165 = 5706.78 m/s^2